OpenStudy (mitu12):

Factor the following polynomials correctly q^2-9q+18

OpenStudy (trojanpoem):

(q - 6)(q-3)

OpenStudy (sweetburger):

what 2 values multiply to equal 18 and add to equal -9... alright nm

OpenStudy (mitu12):

could you explain how to do this

OpenStudy (anonymous):

Your first two automatics are the q's

OpenStudy (anonymous):

so we have (q+-?)(q+-?)

OpenStudy (mitu12):

ok

OpenStudy (trojanpoem):

@sweetburger , that's how I do it. You could complete the square.

OpenStudy (anonymous):

Next thing you have to do is find what multiplies to =18

OpenStudy (anonymous):

so what are some factors of 18

OpenStudy (mitu12):

-6 and -3

OpenStudy (sweetburger):

something interesting to note when the notation is like this ax^2-bx+c the final form will look like this (q+-?)(q+-?). if the form looks like ax^2-bx-c then the notation will look like this (q+-?)(q+?). And finally if it looks like this ax^2+bx+c the notation will look like this (q+?)(q+?).

OpenStudy (anonymous):

correct and then once you get your factors plug them in for the ?'s and test and see if they work. To test it you simply multiply the sets in the parenthesis by each other.

OpenStudy (sweetburger):

@TrojanPoem I do actually know how to do these kind of problems so you dont have to tell me...

OpenStudy (trojanpoem):

@manutdd30 , Well, now he knows them. But how about this ? i.e: 2x^2 - 18 x + 40 @mitu12

OpenStudy (trojanpoem):

@sweetburger , I was telling you I do it the same. BUt complete the square is for mitu12

OpenStudy (trojanpoem):

Work up. your brian.

OpenStudy (sweetburger):

The question in your example you just apply the quadratic or use the rational root theorem and then use the remainder theorem

OpenStudy (trojanpoem):

-,- , Not for you @sweetburger , it's for @mitu12.

OpenStudy (anonymous):

Trying to factor by splitting the middle term 2.1 Factoring q^2-9q+18 The first term is, q^2 its coefficient is 1 . The middle term is, -9q its coefficient is -9 . The last term, "the constant", is +18 Step-1 : Multiply the coefficient of the first term by the constant 1 • 18 = 18 Step-2 : Find two factors of 18 whose sum equals the coefficient of the middle term, which is -9 . -18 + -1 = -19 -9 + -2 = -11 -6 + -3 = -9 That's it

OpenStudy (anonymous):

^ What do you think you should do next?

OpenStudy (mitu12):

the two factors are -6 and -3

OpenStudy (anonymous):

Yes… then what?

OpenStudy (mitu12):

then i think you plug them into the parenthesis

OpenStudy (anonymous):

lol no… Please don't give out direct answers :) @TrojanPoem Instead, try engaging the asker so they can learn better and understand where your answer is coming from. As the saying goes… \(\\\color{darkblue}{Give~a~man~a~fish,~he~eats~for~a~day.}\) \(\\\color{pink}{Teach~a~man~to~fish,~he~eats~for~a~lifetime.}\)

OpenStudy (anonymous):

Step-3 : Rewrite the polynomial splitting the middle term using the two factors found in step 2 above, -6 and -3 q^2 - 6q - 3q - 18

OpenStudy (sweetburger):

Technically there are 3 more factors that should be included which are 18,1 2,9 and 3,6 due to the c value being positive.

OpenStudy (sweetburger):

I know you were limiting them for clarification but just putting that out there.

OpenStudy (anonymous):

Thanks :)

OpenStudy (sweetburger):

Nice job @Aureyliant :).

OpenStudy (anonymous):

@mitu12 What do you believe you should do for step 4? Recall factoring by grouping.

OpenStudy (mitu12):

yes and you will get q(q-6)-3(q-6)

OpenStudy (anonymous):

q2 - 6q - 3q - 18 Step-4 : Add up the first 2 terms, pulling out like factors : q • (q-6) Add up the last 2 terms, pulling out common factors : 3 • (q-6) Step-5 : Add up the four terms of step 4 : (q-3) • (q-6)

OpenStudy (mitu12):

then you will get (q-3)(q-6)

OpenStudy (anonymous):

Yes.. good job.

OpenStudy (mitu12):

thank you

OpenStudy (mitu12):

same method i use just forgot